OpenStudy (anonymous):

A group of college students is volunteering for Help the Homeless during their spring break. They are putting the finishing touches on a house they built. Working alone, Nita can paint a certain room in 3 hours. Jorge can paint the same room in 4 hours. Identify the equation that can be used to find how long it will take them working together to paint the room. How many hours will it take them to paint the room?

OpenStudy (anonymous):

\[\frac{3\times 4}{3+4}\] is what you are looking for

OpenStudy (anonymous):

i didn't know they had that many homeless in ft lauderdale

OpenStudy (anonymous):

we can work through it slowly if you like , but if you do more than two of these you can see how to get \[\frac{3\times 4}{3+4}\] quickly your choice

OpenStudy (shamim):

nita can paint the room in 3 hours

OpenStudy (shamim):

nita can paint in 1 hour 1/3 of the room